OpenStudy (anonymous):

Harry had $32. He spent all the money on buying 3 notebooks for $x each and 4 packs of index cards for $y each. If Harry had bought 5 notebooks and 5 packs of index cards, he would have run short of $18. A student concluded that the price of each notebook is $5 and the price of each pack of index cards is $1. Which statement best justifies whether the student's conclusion is correct or incorrect?

OpenStudy (anonymous):

The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 50 is (8, 2). The student's conclusion is incorrect because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (8, 2). The student's conclusion is correct because the solution to the system of equations 3x + 4y = 32 and 5x + 5y = 18 is (5, 1). The student's conclusion is correct because the solution to the system of equations 3x – 4y = 32 and 5x – 5y = 50 is (5, 1).

OpenStudy (anonymous):

\[3x+4y=32\] is one of the equations then since \(32+18=50\) the other would be \[5x+5y=50\]

OpenStudy (anonymous):

you know how to solve \[3x+4y=32\\ 5x+5y=50\]?

OpenStudy (anonymous):

"no" is a fine answer, just asking i can show you if you like

OpenStudy (anonymous):

no... :/

OpenStudy (anonymous):

DO U STILL NEED HELP

OpenStudy (anonymous):

yes whats the answer

OpenStudy (anonymous):

@hype.child

OpenStudy (anonymous):

its A @Jasmiinnee_m